OpenStudy (anonymous):

Find the exact value by using a half-angle identity. tan(7pi/8)

OpenStudy (anonymous):

-0.414213562373095048801688724209698078569671875376948073176679737990732478462107038850387534327641572735013846230912297024924836055850737212644121497099935831413222665927505592755799950501152782060571470109559971605970274534596862014728517418640889198609552329230484308714321450839762603627995251407989687253396546331808829640620615258352395054745750287759961729835575220337531857011354374603408498847160386

OpenStudy (anonymous):

what is this?

OpenStudy (anonymous):

=sin(7pi/4)/1+cos(7pi/4) =(-1/radical 2)/(1+1/radical2) =-1

OpenStudy (anonymous):

HAHAH mine is more exact :P

OpenStudy (anonymous):

hmm.. ya maybe

OpenStudy (anonymous):

thanks guys .

OpenStudy (anonymous):

I medal'd you solane. Yours was more like what they needed :P

OpenStudy (anonymous):

no prob :) and yahh i think u're answer was a bit too long..just saying :P

OpenStudy (anonymous):

Your right. :)

OpenStudy (anonymous):

what about this :-Rewrite with only sin x and cos x. sin 2x – cos 2x

OpenStudy (anonymous):

sin2x=2sinxcosx

OpenStudy (anonymous):

that's the double angle identity